OpenStudy (anonymous):

Pipe A can fill a tank in 3 hours. If pipe b can fill the same tank in 2 hours, how many minutes will it takes both pipes to fill 2/3 of the tank?

OpenStudy (anonymous):

one fills \(\frac{1}{3}\) of the tank per hour, the other fills \(\frac{1}{2}\) of it their combined rate is \(\frac{1}{3}+\frac{1}{2}=\frac{5}{6}\) per hour solve \[\frac{5}{6}t=\frac{2}{3}\] for \(t\) in one step